To recharge a 10 V battery, a battery charger must move 3.7 105 C of charge from the negative terminal to the positive terminal. How much work is done by the charger? Express your answer in joules.
Answer #1
W =qV = 10*3.7*10^5 = 3.7*10^6 J
Answer #2
work done = charge * potential difference
W = Q * V
=10*3.7*10^5 = 3.7 MegaJoules
Answer #3

Since ΔV=W/q
therefore W = ΔV*q.
= 10*3.7*105
= 37*105 J.

ALL THE BEST :):):)